OpenStudy (anonymous):

The circumference of a circle is 30 millimeters. What is its diameter? Estimate, using 3 for π.

OpenStudy (anonymous):

Please Help!

OpenStudy (asnaseer):

do you know how the circumference of a circle is realted to its radius/diameter?

OpenStudy (anonymous):

not really sorry ;(

OpenStudy (asnaseer):

the relation is as follows:\[C=2\pi r=\pi D\]where C is the circumference, r is the radius and D is the diameter. also note that:\[D=2r\]can you solve it now?

OpenStudy (anonymous):

no don't get it

OpenStudy (anonymous):

can you work it out?

OpenStudy (anonymous):

please

OpenStudy (asnaseer):

do you want to learn how to do this or do you just want the answer?

OpenStudy (anonymous):

i want both please

OpenStudy (anonymous):

are you still here?

OpenStudy (asnaseer):

ok, so we are given the circumference (C=30 mm) and we are asked for the diameter D. so we can use this equation:\[C=\pi D\]we can re-arrange this as:\[D=C/\pi\]we are also told to estimate the answer by using the value 3 for PI, so we get:\[D\approx C/3=30/3=10\]so D is approximately 10 mm.

OpenStudy (anonymous):

thanks alot!